OpenStudy (anonymous):

How many solutions does this system have? -3x + 6y = 10 -3x + 6y = -4

Directrix (directrix):

-3x + 6y = 10-------------------- -3x + 6y = 10 -3x + 6y = -4 ------Multiply by -1 -- 3x - 6y = 4 @trentren Add the like terms. What do you get?

Directrix (directrix):

Here is a chart of the possible relationships.

OpenStudy (anonymous):

what? but there aren't any like terms in -3x + 6 = 10

Directrix (directrix):

Add these two: -3x + 6y = 10 3x - 6y = 4 --------------- The x terms are like terms, the y terms are like terms. The 10 and the 4 are like terms. So just add them and then we can knock out this problem.

OpenStudy (benlindquist):

yea

Directrix (directrix):

Waiting for you to add.

Directrix (directrix):

Okay

OpenStudy (anonymous):

OK so I add 10 + 4? Thats 14

Directrix (directrix):

Correct. What is -3x + 3x =

OpenStudy (anonymous):

0

Directrix (directrix):

Correct

Directrix (directrix):

What is 6x - 6x =

OpenStudy (anonymous):

okay....so

OpenStudy (anonymous):

no solutions?

Directrix (directrix):

-3x + 6y = 10 3x - 6y = 4 --------------- 0 + 0 = 14 0 = 14 No Solution.

OpenStudy (anonymous):

Okay thanks.

Directrix (directrix):

You are welcome.